OpenStudy (anonymous):

4 4/14 + 2 3/7.. can u please show me the steps..

OpenStudy (anonymous):

47/7

OpenStudy (anonymous):

can u show me how u did it? i need to learn.. thanks

OpenStudy (anonymous):

Change it to improper fraction, then make the denominator's equal. Then add

OpenStudy (anonymous):

i m in 5th grade but dont know how.. can u show me?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4 4/14 + 2 3/7 3/7 * 2 = 6/14 4/14 + 6/14 = 10/14 10/14 divided by 2 = 5/7 = 6 5/7 i think thats right, im not too good at math.

OpenStudy (anonymous):